5926000239:

Which line most clearly creates a conflict between the protagonist and the setting? Kayden felt the branches and briars pulling at her, slowing her down, blocking her path. Kayden pushed aside the prickly thorns and moved quietly down the trail, crying as she went. Kayden kept going, despite the pain and guilt she felt inside and the thick brush in her way. Kayden wished the path were shorter, the way easier to walk, and her decision less difficult.

5926000239:

there

G3tH3Pxmpthat5Amx5t:

2+2=4-1=3=C i think.

5926000239:

what the

Mathew222:

yea

Mathew222:

its A

5926000239:

ok thanks